OpenStudy (anonymous):

WILL GIVE METAL FOR THE BEST ANSWER Simplify: square root of 3 times square root of 21 square root of 24 square root of 63 3 square root of 7

OpenStudy (anonymous):

\[\sqrt{3}\sqrt{21} = \sqrt{3}\sqrt{3*7}=\sqrt{3}\sqrt{3}\sqrt{7}=3\sqrt{7}\]

OpenStudy (anonymous):

It is the second chouice!

OpenStudy (anonymous):

choice** @Torri_

OpenStudy (anonymous):

Thank you so much @Tom_Boy_Rebel enjoy your metal!

OpenStudy (anonymous):

Last choice

OpenStudy (anonymous):

second choice isn't simplified...

OpenStudy (anonymous):

Thank you @Torri_ Have a nice day!

OpenStudy (anonymous):

sorry @Tom_Boy_Rebel ... @mtbender74 said that the last choice was the right answer and i agree, i will have to take back the metal and give it to @mtbender74

OpenStudy (anonymous):

thanks @Torri_

OpenStudy (anonymous):

No its fine @Torri_ :D but that means that I need help so can you please explain @mtbender74

OpenStudy (anonymous):

Certainly... What you did was right... \[\sqrt{3}\sqrt{21} = \sqrt{63}\] But to simplify, you need to factor the number and see if you have a pair that can pull out.. \[\sqrt{63}=\sqrt{3*3*7}\] Since you have a pair of 3's, they can come outside of the radical as a single 3...leaving you with \[3\sqrt{7}\] What i did is the same, i just factored the 21 first and then combined the two root 3's...